OpenStudy (anonymous):

What is the eccentricity of the graph of (x-2)^2/121+(y+6)^2/64=1 ? –1.24 –0.68 0.68 1.24

OpenStudy (anonymous):

\[\frac{(x-2)^2}{11^2}+\frac{(y+6)^2}{8^2}=1\] a=11, b=8 \[a^2+b^2=c^2\] \[e=\frac{c}{a}\] You gt it :)